The AM of 2 nos exceeds their GM by 15 & HM by 27 .Find the nos .

Grade:11

1 Answers

Harsh Patodia IIT Roorkee
9 years ago
Let two numbers be x and y
AM = (a+b)/2
HM= (2ab)/(a+b)
GM= \sqrt{ab}

(a+b)/2 = \sqrt{ab} +15
(a+b)/2= 2ab/(a+b) + 27

Solving above you will get 120 and 30 as 2 numbers